OpenStudy (anonymous):

Find the solutions of the equation Cos2(theta)=3Cos(theta)-2 that are in the interval [0,2pi)

OpenStudy (anonymous):

is it \[\cos 2 \theta or \cos^2\theta\]

OpenStudy (anonymous):

cos^2 -3cos +2=0 c0s^2 -2cos -cos+2=0 cos(cos-2)-1(cos-2)=0 (cos-1)(cos-2)=0

OpenStudy (anonymous):

cos=2 not possible as it give max value =1

OpenStudy (anonymous):

cos theta =1

OpenStudy (anonymous):

it give value =1 at 0 and 2pi